OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p) solve for p please

OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p) –6p + 7 = 6p – 9 + 40 - 16p -6p + 7 = -10p + 31 -6p + 10p + 7 = 31 4p + 7= 31 4p = 31 - 7 4p =24 p = 24/4 p = 6